(7.6 #2)
Can someone help me with this?
By what should both sides of the equation below be multiplied to clear it of fractions?
a) x/7 = 2x/3 + 9/4x
b) 6x/x-6 = 5+ 1/x-6
Thank you,
Alberto

Answer by checkley77(12844)

a) x/7 = 2x/3 + 9x/4
x/7=2x/3+9x/4 the LCD for these 3 fractions is 7*3*4=84.
84x/7=84*2x/3+84*9x/4
12x=56x+189x
12X-56X-189X=0
-233X=0
X=0/-233
X=0 ANS.
----------------------------------------
b) 6x/(x-6) = 5+ 1/(x-6) hERE THE LCD=(X-6)
6X(X-6)/(X-6)=5(X-6)+1(X-6)/(X-6)
6X=5X-30+1
6X-5X=-29
X=-29 ANS.